Find the inverse of y=2√(x+3). Thanks!
Answer by lwsshak3(11628)

Find the inverse of y=2√(x+3).
interchange x and y then solve for y
x=2√(y+3)
square both sides
x^2=4(y+3)
x^2=4y+12
4y=x^2-12
y^-1=(x^2-12)/4